The stopping distance, d, of a particular car after the brakes are applied varies directly as the square of the rate, r. If the car is travelling 40 mph, it can stop in 80 feet. How many feet will it take the same car to stop when it is travelling 80 mph?
Answer by Boreal(15235)

d=kr^2
80=k^40^2, solve for k
80/1600=k=0.05
use k in second part
d=k*80^2=0.05*6400
d=320 feet